OpenStudy (anonymous):

Where does the graph of y = 2x2 − 9x − 5 cross the x-axis? (−1 over 2, 0) and (5, 0) (1 over 2, 0) and (−5, 0) (−5 over 2, 0) and (1, 0) (5 over 2, 0) and (−1, 0)

OpenStudy (agent0smith):

\[\Large 2x^2 − 9x − 5=0\] can you factor this...?

OpenStudy (anonymous):

no?

OpenStudy (agent0smith):

Do you know how to factor something like x^2 + 3x + 2 into (x+2)(x+1)?

OpenStudy (anonymous):

yes

OpenStudy (agent0smith):

Okay, well we can use a similar method here... but you have to change it a bit. \[\Large 2x^2 − 9x − 5=0\] Start by multiplying the first number by the last number: 2*(-5) = -10 now you need to find two numbers that multiply to -10, and add up to the middle number, -9...

OpenStudy (anonymous):

10 & -1?

OpenStudy (agent0smith):

Good... but it should be -10 and +1, since they add up to -9 now you need to break up that -9x into -10x and +1x: \[\Large 2x^2 − 10x + x − 5=0\] I'll put thebrackets here for the next step \[\Large ( 2x^2 − 10x) + (x − 5)=0\] now you can factor the 2x out of the first brackets, \[\Large 2x(x − 5) + 1(x − 5)=0\] and one more step...

OpenStudy (agent0smith):

Can you take out the common factor?

OpenStudy (anonymous):

you lost me

OpenStudy (agent0smith):

which step...?

OpenStudy (anonymous):

last one

OpenStudy (agent0smith):

\[\Large ( 2x^2 − 10x) + (x − 5)=0\] notice you can factor a 2x out of the first brackets? it's a common factor of bother 2x^2 and -10x

OpenStudy (anonymous):

ok so how do i get the final answer

OpenStudy (agent0smith):

What common factor do you see here?\[ \Large 2x(x − 5) + 1(x − 5)=0 \] they both have (x-5), so pull that out the front

OpenStudy (agent0smith):

like so (x-5)(2x+1) = 0 now you can easily find the solutions.

OpenStudy (anonymous):

5, -1/2?

OpenStudy (anonymous):

but theres two answers?

OpenStudy (anonymous):

oh wait, nvm, its A?

OpenStudy (agent0smith):

Good job :)

OpenStudy (anonymous):

could you help me with one more?

OpenStudy (agent0smith):

Sure

OpenStudy (anonymous):

Identify the solutions of 2x2 + 13x + 6 = 0. x = 1 over 2, x = 6 x = −1 over 2, x = −6 x = −3 over 2, x = −2 x = 3 over 2, x = 2

OpenStudy (agent0smith):

2x^2 + 13x + 6 = 0 do it the same way. Multiply first by last: 2*6 = 12 now find two factors of 12 that add up to the middle 13.

OpenStudy (anonymous):

12 & 1

OpenStudy (anonymous):

so 2x(x + 12)( x+1) ?

OpenStudy (agent0smith):

2x^2 + 12x + 1x + 6 = 0 (2x^2 + 12x) + (1x + 6) = 0 so factor a 2x out of the first brackets 2x(x+6) + 1(x+6) = 0

OpenStudy (agent0smith):

Then just take out the common factor (x+6)

OpenStudy (anonymous):

x = 6

OpenStudy (anonymous):

and the other one idk

OpenStudy (anonymous):

2x(x+6) <- dont get how to do that

OpenStudy (agent0smith):

hold on, first go from this 2x (x+6) + 1 (x+6) = 0 take out the (x+6) to get: (x+6) (2x+1) = 0 now find your x's. 6 isn't one.

OpenStudy (anonymous):

-6 , -1/2 ?

OpenStudy (agent0smith):

Good :D

OpenStudy (anonymous):

thanks!

OpenStudy (agent0smith):

Welcome :)